Find and simplify f(a+h)-f(a)/h (that is all over h by the way not just f(a) but the f(a+h) part too. (h cannot be 0) for the function f(x)=2-8x^2
thank you so much!

Answer by stanbon(75887)

Find and simplify f(a+h)-f(a)/h (that is all over h by the way not just f(a) but the f(a+h) part too. (h cannot be 0) for the function f(x)=2-8x^2
---------
f(a+h) = 2 - 8(a+h)^2 = 2-8(a^2+2ah++h^2)
= 2 -8a^2 -16ah -9h^2
-------------
f(a) = 2 -8a^2
----------------
Subtracting f(a) from f(a+h) you get;
-16ah - h^2
-----------------
Dividing that result by h you get:
-16a -h
==========================
Cheers,
StanH.